# Qwen vs. Gemma Integration with Whisper: A Comparative Study in Multilingual SpeechLLM Systems

### v1

@@ 1 Introduction
In our submission to the MLC-SLM Challenge [^1] https://www.nexdata.ai/competition/mlc-slm, we propose a streamlined and effective system architecture that harnesses the strengths of pretrained models with minimal task-specific engineering. Our system utilizes OpenAI’s Whisper model [7] as the speech encoder due to its strong generalization capabilities and robustness to multilingual input. For the language modeling component, we explore both Qwen2.5 [8] and Gemma3 [9]. A lightweight linear projector module is trained to bridge the speech and language modalities. Through this simple yet effective setup, we demonstrate competitive performance in multilingual conversational speech modeling, highlighting the strength of modular design and pre-trained components over heavily customized architectures.
@@ 4.1 Main results
Table 2: WER/CER (%) for each language on the development set of the baseline systems and our models. Bold indicates the best result overall (row-wise), and underline indicates the best result within each model group for that language.

### v2

@@ 1 Introduction
In our submission to the MLC-SLM Challenge [^1] https://www.nexdata.ai/competition/mlc-slm, we propose a streamlined and effective system architecture that harnesses the strengths of pretrained models with minimal task-specific engineering. Our system utilizes OpenAI’s Whisper model [7] as the speech encoder due to its strong generalization capabilities and robustness to multilingual input. For the language modeling component, we explore both Qwen2.5 [8] and Gemma3 [9]. A lightweight linear projector module is trained to bridge the speech and language modalities. Through this simple yet effective setup, we demonstrate competitive performance in multilingual conversational speech modeling, highlighting the strength of modular design and pre-trained components over heavily customized architectures [^2] Implementations at: https://github.com/tuanio/i2r-speechllm-mlc-slm.
@@ 4.1 Main results
Table 2: WER/CER (%) for each language on the development set of the baseline systems and our models. Bold indicates the best result overall (row-wise), and underline indicates the best result of every model for each language group.